OpenStudy (anonymous):

If w(t) = 3t + 5, solve w(t - 1) = w(2t)

OpenStudy (anonymous):

w(t-1) = w(2t) 3(t-1) + 5 = 3(2t) + 5 3t - 3 + 5 = 6t + 5 3t - 3 = 6t -3 = 3t -1 = t t = -1